# Cosmic Strings from Tribrid Inflation

### v1

@@ 1 Introduction
In addition to solving the horizon and flatness problems, inflation has the further benefit that it can dilute away unwanted topological defects [5, 3] such as monopoles. In general, monopoles form at some stage of symmetry breaking in Grand Unified Theories (GUTs), because any GUT has to break in such a way that $U(1)_{\text{em}}$ is left unbroken. However, observations give strong upper bounds on the monopole abundance, which is inconsistent with the GUT predictions [6]. This so-called monopole problem can be solved if the GUT breaking generates the monopoles before inflation and if the subsequent period of accelerated expansion lasts long enough to sufficiently dilute the monopole abundance. On the other hand, topological defects such as (metastable) cosmic strings (see e.g. [7] for a review) could form after inflation without causing a similar problem as the monopoles. In fact, if they form after inflation they can even leave an observable signature in the stochastic gravitational-wave background (SGWB) that can provide a valuable window into the early universe and and underlying particle physics models at high energies.
@@ 2.2 Topological Defect Formation
Cosmic strings are one-dimensional topological defects which are formed when the vacuum manifold $\mathcal{M}$ contains unshrinkable loops. This is characterised by the non-triviality of the first homotopy group. When $\pi_{1}(\mathcal{M})\neq[I]$, it means that there exists a loop $f:S^{1}\rightarrow\mathcal{M}$ which cannot be continuously deformed to the the identity map that sends all points from the circle $S^{1}$ to the base point $x\in\mathcal{M}$, i.e. $f$ is a loop which is unshrinkable to a point. Since
@@ 3.5 Cases 1, 2, 3 with $\langle S\rangle\neq 0$
Considering $\langle S\rangle\neq 0$ for case 1, we find that it does not qualitatively change the results for topological defect formation. This is because, although the squared masses of the waterfall field directions change due to the non-zero $S$ vev, we still have $m_{1}^{2}=m_{2}^{2}$ and $m_{3}^{2}=m_{4}^{2}$, where $m_{1,2}^{2}$ are always positive and $m_{3,4}^{2}$ become negative as soon as the critical value is reached (which now also depends on on the value of $S$). This means that there is again only one critical value and at this value two waterfall directions become tachyonic simultaneously. Hence, again cosmic strings form.
@@ 3.6.1 Case 1a
$\langle\bar{\phi}\rangle=0,\quad\langle\phi\rangle=-\frac{\delta\sqrt{\frac{|\zeta|}{\Lambda}}}{\frac{\zeta}{\Lambda}}\frac{1}{\left(M^{4}\frac{|\zeta|^{2}}{\Lambda^{2}}+|\delta|^{2}\right)^{1/4}}e^{i\theta},$ (3.40)
@@ 3.6.3 Case 1c
We note that beyond the examples shown above (cases 1a, 1b and 1c) one could also consider more suppressed linear terms like $\frac{\delta}{\Lambda^{2}}S(\phi\bar{\phi})(\phi\bar{H})$, motivated for instance by the desire to explain the smallness of the deformation from a suppressed effective operator. As long as the deformation results in non-zero $\langle H\rangle,\langle\bar{H}\rangle$ and a unique field trajectory that that leads to the fields ending up in the same vacuum everywhere in the observable universe, the conclusion on the absence of topological defect production remains unchanged.
@@ 3.6.4 Case 1d
However, in contrast to the discussion there, the cubic term deforms the potential such that already before the critical point is reached, another local minimum arises in the potential, which is energetically lower than the one at $\langle H\rangle=\langle\bar{H}\rangle=0$. Above the critical point, however, there is a potential barrier separating the two minima. At the critical point, the potential barrier vanishes and the point $\langle H\rangle=\langle\bar{H}\rangle=0$ becomes a saddle point. From the unstable saddle point, and even a bit earlier by tunneling, (quantum) fluctuations of the waterfall fields can allow them to move to the energetically favoured minimum. On the other hand, below the critical point, where $m^{2}_{3,4}<0$, $\langle H\rangle=\langle\bar{H}\rangle=0$ becomes a maximum and in regions where the waterfall fields are located at this point they can move in any direction of the two-dimensional field space. Compared to case 1, as long as $|\langle\phi\rangle|=|\langle\bar{\phi}\rangle|\not=0$ the potential is tilted, such that there is one unique minimum (which corresponds to the one that formed already before the critical point). When $\langle\phi\rangle,\langle\bar{\phi}\rangle$ roll towards zero, the deformation gets smaller and finally vanishes for $\langle\phi\rangle=\langle\bar{\phi}\rangle=0$.
@@
Effect on Topological Defect Formation: When the critical point is reached and the waterfall field potential features a saddle point, in most regions of the universe the waterfall field will move to the new energetically favoured minimum. However, it is possible that in some smaller part of the universe, the waterfall fields remain at zero field values. From this maximum, they can roll in any direction of the two-dimensional field space, leading to the formation of cosmic strings. Due to the strong bias towards the energetically favoured minimum, we expect that this happens in smaller regions of the universe, such that string loops of reduced size are produced. Furthermore, since the potential is tilted (as long as $|\langle\phi\rangle|=|\langle\bar{\phi}\rangle|\not=0$), on one side of the cosmic strings the potential energy is higher than on the other side, which implies that the string loops form boundaries of temporary domain walls.
@@
It is energetically favoured to reduce the size of these domain walls, which accelerates the shrinking of the cosmic string loops. Depending on various parameters, such as the size of the cubic deformation term and the details of the inflaton potential, the unstable system of domain walls bounded by strings can effectively disappear already before the inflaton fields settle at zero. On the other hand, it might also happen that a reduced number of cosmic string loops survive until this point (where the deformation and thus the domain walls vanish since the potential is not tilted anymore). A more detailed dedicated study of this case, which would involve a careful simulation of the quantum fluctuations of the waterfall fields, is beyond the scope of the present work but would be highly desirable.

### v2

@@ 1 Introduction
In addition to solving the horizon and flatness problems, inflation has the further benefit that it can dilute away unwanted topological defects [5, 3] such as monopoles. In general, monopoles form at some stage of symmetry breaking in Grand Unified Theories (GUTs), because any GUT has to break in such a way that $U(1)_{\text{em}}$ is left unbroken. However, observations give strong upper bounds on the monopole abundance, which is inconsistent with the GUT predictions [6]. This so-called monopole problem can be solved if the GUT breaking generates the monopoles before inflation and if the subsequent period of accelerated expansion lasts long enough to sufficiently dilute the monopole abundance. On the other hand, topological defects such as (metastable) cosmic strings (see e.g. [7] for a review) could form after inflation without causing a similar problem as the monopoles. In fact, if they form after inflation they can even leave an observable signature in the stochastic gravitational-wave background (SGWB) that can provide a valuable window into the early universe and underlying particle physics models at high energies.
@@ 2.2 Topological Defect Formation
Cosmic strings are one-dimensional topological defects which are formed when the vacuum manifold $\mathcal{M}$ contains unshrinkable loops. This is characterised by the non-triviality of the first homotopy group. When $\pi_{1}(\mathcal{M})\neq[I]$, it means that there exists a loop $f:S^{1}\rightarrow\mathcal{M}$ which cannot be continuously deformed to the identity map that sends all points from the circle $S^{1}$ to the base point $x\in\mathcal{M}$, i.e. $f$ is a loop which is unshrinkable to a point. Since
@@ 3.5 Cases 1, 2, 3 with $\langle S\rangle\neq 0$
Considering $\langle S\rangle\neq 0$ for case 1, we find that it does not qualitatively change the results for topological defect formation. This is because, although the squared masses of the waterfall field directions change due to the non-zero $S$ vev, we still have $m_{1}^{2}=m_{2}^{2}$ and $m_{3}^{2}=m_{4}^{2}$, where $m_{1,2}^{2}$ are always positive and $m_{3,4}^{2}$ become negative as soon as the critical value is reached (which now also depends on the value of $S$). This means that there is again only one critical value and at this value two waterfall directions become tachyonic simultaneously. Hence, again cosmic strings form.
@@ 3.6.1 Case 1a
$\langle\bar{\phi}\rangle=0,\quad\langle\phi\rangle=-\frac{\delta\sqrt{\frac{|\zeta|}{\Lambda}}}{\frac{\zeta}{\Lambda}}\frac{1}{\left(M^{4}\frac{|\zeta|^{2}}{\Lambda^{2}}+|\delta|^{2}\right)^{1/4}}e^{i\theta},$
@@ 3.6.3 Case 1c
We note that beyond the examples shown above (cases 1a, 1b and 1c) one could also consider more suppressed linear terms like $\frac{\delta}{\Lambda^{2}}S(\phi\bar{\phi})(\phi\bar{H})$, motivated for instance by the desire to explain the smallness of the deformation from a suppressed effective operator. As long as the deformation results in non-zero $\langle H\rangle,\langle\bar{H}\rangle$ and a unique field trajectory that leads to the fields ending up in the same vacuum everywhere in the observable universe, the conclusion on the absence of topological defect production remains unchanged.
@@ 3.6.4 Case 1d
However, in contrast to the discussion there, the cubic term deforms the potential such that already before the critical point is reached, another local minimum arises in the potential, which is energetically lower than the one at $\langle H\rangle=\langle\bar{H}\rangle=0$. Above the critical point, however, there is a potential barrier separating the two minima. At the critical point, the potential barrier vanishes and the point $\langle H\rangle=\langle\bar{H}\rangle=0$ becomes a saddle point. From the unstable saddle point, and even a bit earlier by tunneling, (quantum) fluctuations of the waterfall fields can allow them to move to the energetically favoured minimum. On the other hand, below the critical point, where $m^{2}_{3,4}<0$, $\langle H\rangle=\langle\bar{H}\rangle=0$ becomes a maximum and in regions where the waterfall fields are located at this point they can move in any direction of the two-dimensional field space.
Compared to case 1, the potential is tilted, such that there is one unique minimum (which corresponds to the one that formed already before the critical point). The global minimum is reached when
$\langle\phi\rangle=0,\quad\langle\bar{\phi}\rangle=-\frac{\delta}{\zeta}\frac{M}{\left(1+\frac{|\delta|^{2}}{|\zeta|^{2}}\right)^{1/4}}e^{i\theta},$
$\langle H\rangle=M\left(1+\frac{|\delta|^{2}}{|\zeta|^{2}}\right)^{1/4}e^{-i\theta},\quad\langle\bar{H}\rangle=\frac{M}{\left(1+\frac{|\delta|^{2}}{|\zeta|^{2}}\right)^{1/4}}e^{i\theta}.$ (3.58)
In this minimum, we have $V=0$ and SUSY is conserved.
@@
Effect on Topological Defect Formation: When the critical point is reached and the waterfall field potential features a saddle point, in most regions of the universe the waterfall field will move to the new energetically favoured minimum. However, it is possible that in some smaller part of the universe, the waterfall fields remain at zero field values. From this maximum, they can roll in any direction of the two-dimensional field space, leading to the formation of cosmic strings. Due to the strong bias towards the energetically favoured minimum, we expect that this happens in smaller regions of the universe, such that string loops of reduced size are produced. Furthermore, since the potential is tilted, on one side of the cosmic strings the potential energy is higher than on the other side, which implies that the string loops form boundaries of temporary domain walls.
@@
It is energetically favoured to reduce the size of these domain walls, which accelerates the shrinking of the cosmic string loops. Depending on various parameters, such as the size of the cubic deformation term and the details of the inflaton potential, the unstable system of domain walls bounded by strings can effectively disappear. A more detailed dedicated study of this case, which would involve a careful simulation of the quantum fluctuations of the waterfall fields, is beyond the scope of the present work but would be highly desirable.
@@ 5 Summary and Conclusions
Acknowledgments
K.T. acknowledges financial support from the Slovenian Research Agency (research core funding No. P1-0035 and PR-12830).